[Hallc_running] [New Logentry] Hall C RC update: Monday, Jan. 17, 2022

gaskelld at jlab.org gaskelld at jlab.org
Mon Jan 17 08:15:01 EST 2022

Logentry Text:
Very productive running for the last 24 hours.  We rotated the SHMS to 6.385 degrees early on day shift yesterday and have been taking data for the SHMS center setting (Q2=3.85, W=3.07, low epsilon) on LD2 since then. At the moment we have 9579 good e-pi events. After one more run, we should be ready to switch to the aluminum dummy target.

Near the end of owl shift, HMS Q1 tripped off - an access will be required to address the problem.

There will be no RC meeting this morning - we will meet again on Tuesday.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3972992
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220117/f30e39f3/attachment.html>

More information about the Hallc_running mailing list